# InversionView: A General-Purpose Method for Reading Information from Neural Activations

### v2

@@ H.1 InversionView Reveals Information from Un-attended Tokens
H.2 Additional Discussion about Path Patching
@@
Appendix I Automated Interpretability
@@
We further explore whether the process of obtaining the common information from a collection of inputs can be automated by LLMs. We use Claude 3444Version: claude-3-opus-20240229 [3] In preliminary experiments we also try GPT-4 [1] but we find Claude 3 works better in our case. In the prompt given to Claude 3, we first describe the task it needs to perform, the terminology we are using (e.g., F1, F2, etc.), the rules (e.g., the pattern it finds should be applicable to each inputs in the $\epsilon$-preimage), input and output form, and the crucial steps it should follow. In addition, we also provide it with 3 demonstrating examples in conjunction with the correct answers. Each example corresponds to a specific activation site and token position (e.g., $a^{0,0}$, A1). In each example, there are 2-3 specific query inputs, each query input is accompanied with 20 (sometimes less) samples that are inside the $\epsilon$-preimage. Claude 3 needs to find the pattern for each query input, and summarize its findings across several query inputs, which is the information contained in general in that activation. In addition to the content described above (the common part shared between prompts), we give it a questioning example, which is the content we would like it to interpret. The questioning example shares the same form as the demonstrating example, except that it contains 5 query inputs and their corresponding samples. In addition, when two separate interpretations are needed based on different A1 value, we run the generation twice with examples of different A1 value, instead of giving the model a mixture of two cases and resorting to its own capacity.
@@
Appendix J Compute Resources
We ran all experiments on an NVIDIA A100 card. Training time for the decoder is between 4 and 6 hours in each of the three case studies, without exhaustively tuning efficiency of the implementation, which we believe could further speed up training. Generation of $\epsilon$-preimage samples, and patching experiments run more quickly, as they require just one pass over the test data. For the largest model (IOI), generation takes 1–2 hours.

### v3

@@ I.1 InversionView Reveals Information from Un-attended Tokens
I.2 Additional Discussion about Path Patching
@@
I.3 Additional Discussion about DLA
Direct Logit Attribution (DLA) extends the logit lens method to study individual model components. Specifically, it projects the output of a model component (thus an update on the residual stream) into vocabulary space, and interpret the component by inspecting the tokens it promotes or suppresses. This method has gained popularity in recent years [23, 50, 16, 19], especially in the research of interpreting factual recall in language models [24, 56, 11, 55]. However, in this section, we argue that DLA is only suitable for studying model components that directly affect the model’s final output, and is not well-suited for components whose effect is mediated by other components. In the circuits found by path patching [50, 13], we can see many components that do not connect to the final output directly, which suggests a substantial part of their effect on the final predictions is mediated by further components. DLA shows their direct effect, which may even be non-causal “side effects". Intuitively speaking, some information is meant to be read by a downstream component, e.g., the S-Inhibition Heads’ output in IOI circuit is meant to be read by the query matrix of Name Mover Heads, and such information may not necessarily be visible when projecting to the vocabulary space. Dao et al. [15] also point out such limitations of DLA.
Attn Head, position Info by InversionView (See Table 2) Top 30 promoted rate Top 30 promoted & 1st name rate Top 30 suppressed rate Top 30 suppressed & 1st name rate
9.9, END IO name 99.5% 98.8% 0% 0%
7.9, END S name; Position of … 0% 0% 3.2% 3.2%
8.10, END S name; Position of … (most of time) 0% 0% 27.4% 7.5%
0.1, S2 S name 0% 0% 0% 0%
0.10, S2 S name 0% 0% 0% 0%
2.2, S1+1 S name (most of time) 0% 0% 0% 0%
4.11, S1+1 S name; Position of … 0% 0% 0% 0%
Table 4: Applying DLA to the heads in IOI circuit. Except the first row, all heads do not directly connect to final output according to the IOI circuit, the results show DLA cannot decode their information. We do not include those heads in which only position information is encoded. “Top 30 promoted (suppressed) rate" means the fraction of input examples where the expected name (IO name for the first row, S name for other rows) is inside the top 30 tokens promoted (suppressed) by the head’s output. “Top 30 promoted (suppressed) & 1st name rate" means the expected name is not only inside the top 30 promoted (suppressed) tokens, but also the most promoted (suppressed) name among a list of common and single-token names, so it does not count when another name is ranked higher. Note that a name can be associated with two tokens (with and without a space before it), when calculating the rate, either of them satisfying the condition will count. The rate is calculated over 1000 random IOI examples. As we can see, except for the first row, the expected name is not observable most of the time.
In Table 4 we show the results of applying DLA to attention heads’ output in IOI circuit. We can see that the expected information is not visible in most cases. The best result comes from the S-Inhibition head 8.10, with only 7.5% of cases where the expected name is in the top-30 tokens and there is no other name being suppressed more than it. The rare cases where the expected name is visible can also be explained by the small direct effect on the final output as depicted by Figure 3(b) in [50].
Therefore, researchers should be cautious when using DLA and should be aware of its limitations. A good usage example is the IOI circuit [50], where the authors first identify those attention heads directly affecting the final logits, and only apply DLA to them and design other experiments to interpret other components. Importantly, in the context of factual recall task, we find that the information given by InversionView about the upper layer attention heads is often visible via DLA, indicating these heads contribute to model’s output directly. Thus, our results can serve as confirmation that prior results relying on DLA in this task are generally reliable.
Appendix J Automated Interpretability
@@ I.3 Additional Discussion about DLA
We further explore whether the process of obtaining the common information from a collection of inputs can be automated by LLMs. We use Claude 3141414Version: claude-3-opus-20240229 [3] In preliminary experiments we also try GPT-4 [1] but we find Claude 3 works better in our case. In the prompt given to Claude 3, we first describe the task it needs to perform, the terminology we are using (e.g., F1, F2, etc.), the rules (e.g., the pattern it finds should be applicable to each inputs in the $\epsilon$-preimage), input and output form, and the crucial steps it should follow. In addition, we also provide it with 3 demonstrating examples in conjunction with the correct answers. Each example corresponds to a specific activation site and token position (e.g., $a^{0,0}$, A1). In each example, there are 2-3 specific query inputs, each query input is accompanied with 20 (sometimes less) samples that are inside the $\epsilon$-preimage. Claude 3 needs to find the pattern for each query input, and summarize its findings across several query inputs, which is the information contained in general in that activation. In addition to the content described above (the common part shared between prompts), we give it a questioning example, which is the content we would like it to interpret. The questioning example shares the same form as the demonstrating example, except that it contains 5 query inputs and their corresponding samples. In addition, when two separate interpretations are needed based on different A1 value, we run the generation twice with examples of different A1 value, instead of giving the model a mixture of two cases and resorting to its own capacity.
@@
Appendix K Compute Resources
We ran all experiments on NVIDIA A100 cards. The decoder is trained for 4-6 hours on 2 GPUs for the first three case studies, without exhaustively tuning efficiency of the implementation, which we believe could further speed up training. Regarding the factual recall task, we train the decoder for less than 1 day on 4 GPUs. Generation of $\epsilon$-preimage samples (including forward passes on the probed model to calculate distance metric) is fast for the first 3 tasks, and it takes around 9 hours on 4 GPUs for factual recall task (for 200 query inputs). Patching experiments run quickly, as they are done for small models.